You borrow money (take out a mortgage) to buy a house. You borrow $800,000 which you will pay back with 10 equal payments made at the end of each of the next 10 years. The annual interest rate is 7 percent. Your first payment will be ____ principal payment, and _____ interest paid.
Answer #1

Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

800,000=Annuity[1-(1.07)^-10]/0.07

800,000=Annuity*7.023581541

Annuity=800,000/7.023581541

=$113,902.00(Approx)

Hence interest payment=$800,000*7%=$56,000

Hence principal payment=(113902-56000)=$57902(Approx).